# Graph Language Models

### v1

@@ 4 Experiments
4.1 Dataset
We evaluate our method on the task of relation label classification on randomly selected subgraphs from the largest connected component of the English part of ConceptNet (CN) version 5.7 (Speer et al., 2017), which consists of 125,661 concepts and 1,025,802 triplets. We select 17 distinct relation label classes (cf. Appendix, Tab. 4), ensuring sufficient frequency and semantic dissimilarity. For each class, we randomly sample 1,000 triplets, allowing only cases where exactly one triplet connects the head and tail entities, to reduce label ambiguity. These 1,000 instances are split into train (800), dev (100), and test (100). This creates a balanced dataset of 13,600 train instances, 1,700 dev instances, and 1,700 test instances. To predict relation labels, we replace them with <extra_id_0>, T5’s first mask token.
To investigate the impact of varying graph complexities, we experiment with different graph sizes of CN subgraphs. We start with a radius of $r=1$, when we consider only the two concepts (head and tail) used by the target triplet. To create a larger graph context, we randomly select 4 adjacent triplets – 2 for the head, and 2 for the tail entity of the original triplet. A graph with radius $r=2$ is formed by the subgraph spanned by all entities used in these 5 triplets. For $r=3$ we again randomly select 2 triplets for each of the outer (up to) 4 entities, yielding (up to) 13 triplets. To avoid accidentally adding more short-ranged information, we restrict the new triplets to triplets that actually extend the radius of the graph. This enables us to control graph size and complexity, while still enabling sufficient diversity in the graph structure. Further, the graphs are created such that graphs for smaller radii are strict subgraphs of graphs with larger radii. This ensures that performance changes with increasing radii are due to long-ranged connections, and not due to potentially different short-ranged information. Tab. 1 shows statistics of train graphs, depending on their radius.
@@ 4.1 Dataset
The correct class label of the targeted relation mainly depends on the concepts in the original triplet. To evaluate model effectiveness when long-ranged connections are crucial, we experiment with masking complete subgraphs around the relation to be predicted. The size of the masked subgraph is denoted by $m$, where $m=0$ means no masking, $m=1$ masks neighboring concepts, $m=2$ masks neighboring concepts and the next relations, and so on. [^4] Formally, $m$ denotes the radius of the masked graph in Levi representation, which should not be confused with the radius $r$ in normal representation. We replace each concept and relation in the masked subgraph with a different mask token.
4.2 Experimental setup
@@
Input to the models is a CN subgraph, where the relation to be predicted is replaced with <extra_id_0>. The GLMs encode the graph as detailed in §3, producing an embedding for each token. To obtain a prediction for the masked relation label, we train a linear classification head that takes the mask’s embedding as input and predicts a probability for each of the 17 relation label classes. As discussed in §3, we verbalize unmasked relations according to static templates, shown in the Appendix, Tab. 4 (Plenz et al., 2023). This makes individual triplets more similar to natural language.
@@
In a standard finetuning setting we train the GLM parameters and the classification head jointly. However, since the GLM is initialized from a LM, we hypothesize that it should produce meaningful embeddings, even without any training. To test this hypothesis, we train only the classification head, i.e., we only train a linear probe on the embedding we obtain from the encoded graph. Note that in this setting, the GLM was never trained on any graph data, similar to a zero-shot setting. The linear probe can only perform linear feature extraction and hence, can only achieve high performance if the GLM embeddings show expressive features without any training.
@@
We report mean accuracy across 5 different runs. Hyperparameters, including number of trainable parameters, are shown in the Appendix, Tab. 5. Unless stated otherwise, we run our experiments with T5-small to allow for many baselines. We also run experiments with T5-base and T5-large to observe scaling with larger model sizes.
@@
Table 2: Relation label classification scores in accuracy in %. Results are shown for Linear Probing and Finetuning.
4.3 Baselines
We compare to several baselines, inspired by related work. For all baselines we utilize the T5 encoder as underlying LM, to avoid any effects due to differences from diverse backbone models.
@@ 4.3 Baselines
To assess the performance of LM-based approaches we linearize the input graphs by concatenating all verbalized triplets to form a sequence. There are many sophisticated and structured ways to linearize different kinds of graphs (e.g., the Penman notation for AMRs (Banarescu et al., 2013)). However, such graph traversals generally require the graph to be rooted, directed and acyclic – which makes them inapplicable to linearizing GoTs. Instead, we order the triplets either randomly (T5 set) or alphabetically (T5 list). For T5 set, the order of triplets is shuffled randomly in every training epoch such that the model can learn to generalize to unseen orderings. The concatenated triplets are then passed to the T5 encoder, and the embedding of <extra_id_0> is presented to the classification head.
@@
For the GNN baselines we first encode each node in the original graph (cf. Fig. 1(a)) with the T5 encoder, and then train a GNN using these static embeddings. After the final layer, the GNN returns 17 logits for each node. As the final logits, we take the mean logit of the two nodes adjacent to the relation to be predicted.
We experimented with different variants as GNN layers: GCN (Kipf and Welling, 2017) and GAT (Veličković et al., 2018). In preliminary experiments we tested different numbers of layers (2, 3, 4, 5), hidden channel dimensions (32, 64, 128), and non-linearities (ReLU, leaky ReLU). The best hyperparameters are shown in Appendix, Tab. 5.
Since GNNs do not come with pretrained weights, we only consider them in the finetuning setting, where all parameters are trained.
@@
We also compare the GLM to models with the same architecture, but with random weight initialization – i.e., normal graph transformers (GTs). This allows us to investigate the impact of weight initialization from a LM. This gives us two more baselines, the $\bm{\ell}$GT and the $\bm{g}$GT. We only consider GTs in the finetuning setting, where parameters are updated during training.
4.4 Results
@@
Beyond a radius of $r=3$, the LM baselines have decreasing performances with increasing radii. By contrast, both $\ell$GLM and $g$GLM show increasing performances with increasing radii. This indicates that GLMs can utilize the additional context. However, performance of the $\ell$GLM lags behind the $g$GLM starting from $r=2$. The LM baselines, on the other hand don’t have any inbuilt methods to grasp distances in the graph, which causes them to fail at distinguishing relevant from less relevant information. The performance gap between $g$GLM and the LM models tends to increase for larger $m$, i.e., when larger sub-structures are masked. But $\ell$GLM underperforms for large $m$, sometimes being even surpassed by LM baselines. This highlights the advantage of the global view in $g$GLM compared to $\ell$GLM when long-ranged connections are necessary.
The overall high performance confirms our assumption that the GLM is compatible with LM weights, even without any further training. Increasing performance with increasing radii further shows that GLMs have good inductive graph biases. When long-range connections are relevant, the representations learned by $g$GLM outperform the locally constrained $\ell$GLM – which showcases the strength of the global view that the $g$GLM is able to take.
Model $r$ 1 2 3 4 5 4 4 4 4 4
$m$ 0 0 0 0 0 1 2 3 4 5
Table 3: Relation label classification scores in accuracy in % when training all parameters. Best score per model family is underlined, and best score overall is bold faced.
Finetuning
Table 2 shows the results when training all parameters. In this setting, models can adjust to the given task, and learn to reason over graphs through parameter updates. Also, the GLM can tune parameters to better match the new architecture. The GLM and LM variants are consistently better than GNN and GT methods, which indicates that linguistic understanding is potentially more important than graph reasoning for this task. Further, the scores for GLM and LM are better than in the linear probing setting, which shows that finetuning is, as expected, beneficial.
Overall, the GLMs perform best, while the GTs perform the worst. The only difference between the two model groups is weight initialization – GLMs are initialized from T5, while GTs are initialized randomly. This highlights the necessity of proper weight initialization. Given additional training data, GTs might have the potential to catch up. However, we anticipate that weight initialization from a LM will offer advantages even in situations where more graph data is accessible. Further, we observe that for $r\geq 1$ and $m=0$ the local GT ($\ell$GT) significantly outperforms its global counterpart $g$GT. For the GLMs the global version is on par, or even better than the local version. This shows the effectiveness of T5’s attention mechanism: thanks to its weight initialization, $g$GLM is capable of attending to relevant tokens even in large context windows. The $g$GT on the other hand, suffers from potentially distracting long-ranged information.
For $m=0$ the differences between GLM and LM approaches are small, with a slight trend for GLMs to outperform LMs on large graphs, and vice versa for small graphs. However, when graph reasoning is more important due to masking ($m\geq 1$), then GLMs consistently and significantly outperform all other baselines. This shows that LMs can learn to do simple graph reasoning through parameter updates, but underperform in more complex graph reasoning tasks where either graphs are larger, or long-ranged connections are required.
For $m\geq 1$, the $g$GLM outperforms $\ell$GLM due to its global connections. In contrast to the linear probing setting, the $\ell$GLM outperforms other baselines for all non-zero levels of masking. This indicates that $\ell$GLM can learn to use long-ranged information during training, if the task requires it. However, in such cases the $g$GLM should probably be preferred due to its higher performance.
Impact of model size
To investigate the effect of model size, we train the most promising approaches (GLM and LM) in 3 different sizes. Tab. 3 shows that overall larger models perform better, which is in line with most research. Surprisingly, the base models sometimes outperform the larger models for settings that require more graph reasoning, i.e., larger $m$. However, these differences are small and non-significant. In most cases, $g$GLM large or base are the best model.
5 Related Work
LMs
One approach to augmenting LMs with knowledge from KGs (Pan et al., 2024) is to formulate pretraining objectives that operate on the KG. E.g., LMs can be trained to generate parts of a KG, which requires LMs to store the KG’s content in its parameters. Typically, individual triplets are used for pretraining (Bosselut et al., 2019; Wang et al., 2021; Hwang et al., 2021; West et al., 2023). In these cases, the graph structure is not a target of pre-training. Some work focuses on generating larger substructures, such as paths or linearized subgraphs (Wang et al., 2020a; Schmitt et al., 2020; Huguet Cabot and Navigli, 2021). In either case, the LM has to memorize the KG, as it will not be part of the input during inference.
Another approach is to provide the linearized KG as part of the input during inference. This is a common approach for KG-to-text generation (Schmitt et al., 2020; Ribeiro et al., 2021; Li et al., 2021), where models are trained to take the linearized graphs as input. A more recent trend is retrieval augmented generation, where relevant parts of a knowledge base – e.g., a KG – are retrieved, and then linearized such that they can be provided as part of a prompt to provide additional context (Yu et al., 2022) (cf. https://www.llamaindex.ai and https://www.langchain.com).
In both approaches the graph needs to be linearized, as it is either the input or the output of a sequence-to-sequence LM. Hence, no graph priors can be enforced – instead, the LM has to learn the graph structure implicitly. In contrast, the GLM models the graph as a true graph and has inductive graph priors instilled in its architecture. This prepares the GLM for more proficient graph reasoning, compared to classical LM-based approaches.
GNNs
LMs excel at capturing the meaning of individual triplets, but struggle with structural reasoning. To alleviate this problem, LMs can be combined with GNNs. A common approach is to obtain node and edge features from LMs, and to aggregate this information in the graph using GNNs (Lin et al., 2019; Malaviya et al., 2020; Zhao et al., 2023). Orthogonal to these approaches, Galkin et al. (2023) use structural properties of the graph to obtain node and relation features, without relying on any LMs.
While some of these approaches allow joint training for textual understanding and graph reasoning, none of them offer a unified method for both. By contrast, our GLM formulation seamlessly integrates language understanding and graph reasoning in one method, providing a holistic framework for embedding language and KGs.
Graph Transformers
GTs, a special type of GNN, got increasingly popular in NLP and other domains (Min et al., 2022; Müller et al., 2023). Koncel-Kedziorski et al. (2019) and Wang et al. (2020b) train GTs to generate text from KGs and AMRs, respectively. Most relevant to our work is Schmitt et al. (2021), who use GTs for KG-to-text generation. They construct a PE matrix similar to our approach, but train their model from scratch, which limits its general applicability: E.g., their model trained on WebNLG (Gardent et al., 2017) has a vocabulary size of only 2,100. In contrast, we initialize our model from T5, which equips our model with the full 32,128 token vocabulary of T5, without any training.
6 Conclusion
In this work we present a new type of Language Model architecture: the Graph Language Model (GLM). The GLM realizes a graph transformer architecture initialized with weights from a LM. This enables it to perform graph reasoning, while simultaneously encoding textual triplets in the graph. We probe this model in a relation classification task that challenges a model’s graph-structure reasoning capabilities, and assess the impact of the GLM’s LM capacity in a zero-shot setting. The GLM surpasses competing baselines in two distinct scenarios: a linear probing setting, where only a linear classification head is trained, and a fully supervised setting, where all parameters undergo training. We anticipate that the GLM will serve as a valuable tool for advancing research in embedding and leveraging knowledge graphs.
Appendix
The appendix contains matrices $P$ and $M$ for $\ell$GLM and $g$GLM in Fig. 4, the relation verbalizations for ConceptNet in Tab. 4, and Hyperparameters in Tab. 5.
(a) Relative positions $P$ for dog in $\ell$GLM

### v3

@@ 5 Experiments
We assess the GLMs’ capabilities for embedding GoTs in two experiments on relation (label) classification, i.e., classifying which relation belongs to a given head and tail entity. One experiment uses ConceptNet (CN; Speer et al., 2017) subgraphs that we construct to enable analysis of the impact of structural graph properties. In a second experiment on Wikidata (Vrandečić and Krötzsch, 2014) subgraphs and associated Wikipedia abstracts we test GLMs on interleaved inputs of text and graph.
5.1 Representing and reasoning over Graphs
We construct a balanced dataset of English CN subgraphs consisting of 13,600 train, 1,700 dev and 1,700 test instances with 17 distinct relations as labels. We replace the relation to be predicted with <extra_id_0>, T5’s first mask token.
To investigate the impact of varying graph complexities, we experiment with different graph sizes denoted by their radius $r$. We ensure that small graphs are strict subgraphs of larger graphs, such that potential performance gains in larger graphs must stem from additional long-ranged context.
@@ 5.1 Representing and reasoning over Graphs
To evaluate model effectiveness when long-ranged connections are crucial, we mask complete subgraphs around the relation to be predicted. The size of a masked subgraph is $m$, where $m=0$ means no mask, $m=1$ masks neighboring concepts, $m=2$ masks neighboring concepts and the next relations, etc. We replace each masked concept and relation with a different mask token. Construction details and statistics are shown in §B.1.1.
5.1.1 Experimental setup
@@
The input to our model is a CN subgraph. The relation to be predicted is replaced with <extra_id_0>. The GLM encodes the graphs as in §4, producing an embedding for each token. A linear classification head gives the final prediction from the mask’s embedding. We verbalize unmasked relations using static templates (Plenz et al., 2023), shown in §B, Table 4.
@@
In a finetuning setting we train the GLM and the classification head jointly. However, since the GLM is initialized from a LM, we hypothesize that it should produce meaningful embeddings, even without any training. To test this hypothesis, we train only the classification head, i.e., we only train a linear probe. In this setting, the GLM was never trained on any graph data, similar to a zero-shot setting. The linear probe only extracts linear features and hence, can only achieve high performance if the GLM embeddings show expressive features.
@@
We report mean accuracy across 5 different runs. See §B.1.2 for hyperparameters. Unless stated otherwise, we use T5-small to allow many baselines.
@@
Table 1: Relation label classification accuracy on CN in %. Results are shown for Linear Probing and Finetuning.
5.1.2 Baselines
We compare to several baselines inspired by related work. For all baselines we utilize the T5 encoder as underlying LM. This allows us to focus on the architectural design of different model types.
@@ 5.1.2 Baselines
For LM-based approaches we linearize the input graphs to a sequence, by concatenating the verbalized triplets. There are structured ways to linearize graphs, but such graph traversals generally require the graph to be directed and acyclic – which makes them inapplicable to linearizing GoTs. Instead, we order the triplets either randomly (T5 set) or alphabetically (T5 list). For T5 set, triplets are shuffled randomly in every training epoch such that the model can learn to generalize to unseen orderings. The concatenated triplets are passed to the T5 encoder, and the embedding of <extra_id_0> is presented to the classification head.
@@
For GNN baselines we encode each node of the original graph (cf. Fig. 2(a)) with the T5 encoder, and train a GNN using these static embeddings. After the final layer, the GNN returns 17 logits for each node. As final logits, we take the mean logit of the two nodes adjacent to the relation to be predicted. We experiment with different variants as GNN layers: GCN (Kipf and Welling, 2017) and GAT (Veličković et al., 2018). Since GNNs do not come with pretrained weights, we only apply them in finetuning, when training all parameters.
@@
Finally we compare GLMs to models with the same architecture, but random weight initialization (normal graph transformers). This allows us to assess the impact of weight initialization from a LM with two further baselines: $\bm{\ell}$GT and $\bm{g}$GT. We only consider GTs with finetuning.
5.1.3 Results
@@
Tab. 1 shows the relation label prediction accuracy for linear probing, i.e., when training only the classification head. Our first observation is that $g$GLM is consistently the best, outperforming $\ell$GLM and the LM baselines. For a radius of $r=1$ we have exactly one triplet, which has almost the same representation in the GLM and LM approaches. The only difference is that the LM baselines have an end-of-sentence token, which the GLM does not have. Surprisingly, not having the end-of-sentence token seems to be an advantage with linear probing, but we will see later that this changes when updating model weights.
For $r\geq 3$, LM baselines show decreasing performance with increasing radii. By contrast, both $\ell$GLM and $g$GLM show increasing performances with increasing radii. This indicates that GLMs can utilize the additional context. But LM baselines don’t have any inbuilt methods to grasp distances in the graph, which could cause them to fail at distinguishing relevant from less relevant information.
The performance gap between $g$GLM and LM models tends to increase for larger $m$, i.e., when larger sub-structures are masked. However, the $\ell$GLM underperforms for large $m$, highlighting the advantage of the global view in $g$GLM when long-ranged connections are necessary.
The overall high performance of GLMs confirms our assumption that GLMs are compatible with LM weights, even without any training. Increasing performance with increasing radii further shows that GLMs have good inductive graph biases. When long-range connections are relevant, the representations learned by $g$GLM outperform the locally constrained $\ell$GLM – which showcases the strength of the global view that the $g$GLM is able to take.
Finetuning
Tab. 1 shows results when training all parameters. In this setting, models can adjust to the task and learn to reason over graphs through parameter updates. In addition, GLMs can tune parameters to better match the novel input structure.
The GLM and LM variants are consistently better than GNN and GT methods, which indicates that linguistic understanding is potentially more important than graph reasoning for this task. Models outperform their linear probing scores, which shows that finetuning is, as expected, beneficial.
Overall, the GLMs perform best, while GTs perform the worst. The only difference between the two model groups is weight initialization – the GLMs are initialized from T5, while the GTs are randomly initialized. Further, we observe that for $r\geq 1$ and $m=0$ the local GT ($\ell$GT) significantly outperforms its global counterpart $g$GT. For the GLM the global version is on par, or even better than the local one. This shows the effectiveness of T5’s attention mechanism: thanks to its weight initialization, $g$GLM attends to relevant tokens even in large context windows, while $g$GT suffers from potentially distracting long-ranged information.
For $m=0$ the differences between GLM and LM approaches are small, with a slight trend for GLMs to outperform LMs on large graphs, and vice versa for small graphs. However, when graph reasoning is more important due to masking ($m\geq 1$), then GLMs consistently and significantly outperform all other baselines. This indicates that LMs can learn to do simple graph reasoning through parameter updates, but underperform in more complex graph reasoning tasks where either graphs are larger, or long-ranged connections are required.
For $m\geq 1$, the $g$GLM outperforms $\ell$GLM due to its global connections. In contrast to the linear probing setting, the $\ell$GLM outperforms other baselines for all non-zero levels of masking. This indicates that $\ell$GLM can learn to use long-ranged information during training, if the task requires it.
Impact of model size
To investigate the effect of model size, we train the most promising approaches (GLM and LM) in 3 different sizes. Tab. 6 in §B.1.3 shows that overall larger models perform better. Surprisingly, the base models sometimes outperform the larger models for settings that require more graph reasoning, i.e., larger $m$. However, these differences are small and non-significant. In most cases, $g$GLM large or base are the best model.
(a) Relation label classification.
(b) Source classification.
Figure 4: KG population test results during training. $g$GLM outperforms T5 set by up to 6 points in 4(a).
5.2 Jointly representing Graph and Text
We now investigate GLM capabilities to process interleaved inputs of text and graph in a KG population setup, i.e., extending a KG with new relation instances. Subtask 1 performs text-guided relation classification where some relations may be inferrable from the text, while others may exploit graph knowledge to make predictions. In Subtask 2, models classify the source of a predicted relation, i.e., whether it can be inferred from the text, or whether it requires (additional) graph knowledge.
We construct our data from Huguet Cabot and Navigli (2021), who offer a corpus of Wikipedia abstracts that are linked to Wikidata via entity linking. Their focus is relation extraction, so they filter the graphs using NLI, such that all triplets are entailed by the text. We augment the entailed triplets with further triplets from Wikidata that are not entailed by the text. For a given text, subgraph, head and tail entity, models will jointly predict the relation and the source. We adopt the 220 most common relations in our train graphs and a “no-relation” label. For source labels we have 3 classes: entailed by the text, not entailed and no-relation. No-relation is the correct label iff the relation is also no-relation. §B.2.1 shows statistics and construction details.
5.2.1 Experimental setup and baselines
Unlike §5.1.1, models now receive text and graph data as input. We train two distinct classification heads on the mask’s embedding for relation and source classification. While the mask is part of the graph, its embedding depends on both modalities. The final loss is the sum of the relation classification and the source prediction loss, weighted by 0.9 and 0.1. We use T5-large, but otherwise baselines are as in §5.1.2. §B.2.2 shows the training details.
5.2.2 Results
Fig. 4 and Tab. 8 show test set performance for a) relation and b) source classification, at different training stages. $g$GLM performs the best overall, followed by $\ell$GLM. LM baselines are competitive, but lag behind at early stages and for source prediction. Again, GT baselines perform poorly, showcasing the advantage of weight initialization in GLM – even with large-scale training data. For all models, training plateaus beyond $\sim$ 500k seen instances (cf. Fig. 8 in §B.2.3), so we stop training at this cut-off.
Tab. 2 gives results for ablating different input modalities to GLMs. Since source prediction always requires text input, we test relation classification w/o source prediction. Ablating the text or graph lowers performance by similar amounts, indicating that GLMs utilize both modalities. Training curves in Fig. 10 reveal that first, the model almost exclusively utilizes text data, but quickly learns to make use of the graph. For textually entailed triplets, text is more impactful than the graph, and vice versa for other triplets (cf. Tab. 9). Ablating graphs lowers source prediction by $\sim$4.5 points, which shows that GLMs benefit from graph information even for predominantly text oriented tasks.
The results show that GLMs can efficiently reason over interleaved inputs of graph and text, especially with limited training data. This makes GLMs a promising new model type for knowledge-intense NLP tasks, such as KG population or Q&A.
Ablation Relation classification Source classification
$\ell$GLM $g$GLM $\ell$GLM $g$GLM
w/ text & graph 82.63 82.25 83.39 83.21
w/o text -6.22 -5.84
w/o graph -6.05 -5.10 -4.67 -4.49
w/o text & graph -19.62- -19.24-
Table 2: Ablations for KG population in macro F1.
6 Conclusion
We present the Graph Language Model (GLM) – a graph transformer initialized with weights from a LM. It excels at graph reasoning, while simultaneously encoding textual triplets in the graph as LMs do, thereby bridging the gap between LMs and GNNs. GLMs can natively reason over joint inputs from texts and graphs, leveraging and enhancing each modality. Experiments show the GLM’s advantage over LM and GNN based baselines, even in a linear probing setting. In particular, GLMs greatly outperform graph transformers. This highlights the need for pretrained LM weights, even for graph reasoning. We therefore advocate GLMs as a valuable tool for advancing research in embedding and leveraging knowledge graphs for NLP tasks.
Limitations
While GLMs are designed as general purpose tools for knowledge-intense NLP tasks, our evaluation is limited to English knowledge graphs. However, we explore various types of knowledge graphs (commonsense and factual) and tasks (relation classification, text-guided relation classification, and source prediction), broadening our empirical assessment. Confirming GLMs improved text and graph reasoning skills for different languages, domains and tasks is left for future work.
Our GLM framework supports instantiation from any LM with relative positional encoding, including rotary positional encoding. Comprehensive comparisons to determine the most suitable models for the GLM framework remain for future investigation. Nonetheless, bidirectional LMs are expected to perform best in the novel framework, because unidirectional LMs necessitate additional inverse relations, as discussed in §4.
Ethical considerations
We do not foresee immediate ethical concerns for our research, as we rely on well-established datasets. However, even established datasets can contain undesirable biases which our method could potentially spread and amplify.
Looking ahead, our focus lies in enriching knowledge graph integration within language models, with the aim of enhancing factuality and mitigating hallucination. This advancement is expected to bolster the reliability and controllability of LMs, leading to positive societal impacts. Furthermore, LMs relying on knowledge graphs may facilitate easier maintenance, potentially reducing the need for frequent retraining of deployed models, thereby promoting sustainability in NLP practices.
Acknowledgements
@@
This work was funded by DFG, the German Research Foundation, within the project “ACCEPT: Perspectivized Argument Knowledge Graphs for Deliberation”, as part of the priority program “RATIO: Robust Argumentation Machines” (SPP-1999).
Appendix A Model
Fig. 5 shows the matrices $P$ and $M$ for $\ell$GLM and $g$GLM. Fig. 6 shows the same matrices for joint encoding of text and graph data.
(a) Relative positions $P$ for dog in $\ell$GLM.